Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 135     Entry time: 27 Jul 2024, 23:42
Author: Romain Letrun 
Group: SPB 
Subject: Shift summary 

X-ray delivery
- 9.22 and 11.0 keV, 1-352 pulses @ 1.1 MHz, 600-1400 uJ.
- Large intensity gradient with long trains (>70% drop with 352 pulses @ 1.1 MHz).
- Very unstable delivery from 20:30 until 23:00.

Optical laser delivery
- 800 nm, 50 fs, 1-113 pulses @ 564 kHz.
- Stable delivery

Achievements/Observations
- Instrument alignment.
- AGIPD timing scan.
- AGIPD flat field measurement at 1.1 MHz.

Issues
- Twice more time required to collect AGIPD flat fields due to large intensity gradient in the train.
- XTD2 1.2 mm CVD unusable. Initially stuck in the beam, but could be retracted with the help of DOC/PLC OCD.

ELOG V3.1.4-7c3fd00